# Bridging Conversational and Collaborative Signals for Conversational Recommendation

### v1

@@ Abstract.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Recommender systemsccs: Computing methodologies Natural language generationccs: Computing methodologies Learning latent representations
@@ 1. Introduction
Conversational recommender systems (CRS) aim to elicit user preferences and generate recommendations through interactive dialogues (He et al., 2023; Li et al., 2018). Unlike traditional recommenders, CRS leverages conversational context to adapt to user preferences. However, existing CRS approaches face two main challenges: (1) They do not leverage collaborative filtering (CF) signals with conversational context, leading to suboptimal recommendations (Zhou et al., 2020; Friedman et al., 2023), and (2) CRS datasets are typically sparse, making it difficult to develop robust item representations for accurate recommendations. To this end, a typical CRS consists of two components: (1) a generator to produce natural language responses and (2) a recommender to rank items aligned with user preferences (He et al., 2023; Li et al., 2018; Chen et al., 2019).
From a CRS model perspective, CF-based approaches excel at modeling user-item relationships but fail to leverage conversational context (Salem et al., 2014; Kang and McAuley, 2018). Similarly, knowledge-driven methods improve semantic understanding using external knowledge graphs but lack the collaborative signals needed to model item relationships (Zhou et al., 2020). Large language models (LLMs) show promise in CRS due to their natural language understanding capabilities (He et al., 2023), but they rely at most on textual metadata, which constrains their ability to capture item relationships (He et al., 2025; Yang et al., 2024). Recent methods integrate CF embeddings with LLMs to improve recommendation tasks. For instance, A-LLMRec incorporates CF embeddings into prompts, while CoLLM maps them onto token embeddings (Kim et al., 2024; Zhang et al., 2023). ILM aligns CF embeddings with textual embeddings but lacks conversational context (Yang et al., 2024). While these methods incorporate CF with LLMs to improve recommendations, they are not designed to utilize conversational context. Our framework bridges this gap by integrating CF signals and conversational data into a unified CRS approach.
@@
In this work, we make two key contributions: First, we construct a new Reddit-ML32M dataset by linking the Reddit-Movie dataset (He et al., 2023) with the MovieLens 32M dataset (Harper and Konstan, 2015). This dataset enriches the interaction space, increasing the total number of interactions from 51,148 to over 30 million and improving dataset density from 0.013% to 0.68%, as shown in Table 1. The dataset facilitates the development of richer item representations for CRS tasks by exploiting both conversational data with CF information at the item level. Second, we propose a framework that integrates conversational context from LLMs with CF-based item representations as depicted in Figure 1. GPT-3.5t is used in a zero-shot setting as a backbone to generate initial recommendations from conversational context, which are then refined by ranking them against item embeddings generated by a pre-trained SASRec model (Kang and McAuley, 2018). This approach mitigates biases in LLM outputs, such as over-reliance on popular items or training data (He et al., 2023; Yang et al., 2024), while incorporating collaborative signals. Empirical evaluation across three paradigms: CF-based recommenders, CRS models, and LLM-based CRS without CF signals shows the effectiveness of our framework. Our approach achieves a 12.32% improvement in Hit Rate and a 9.9% improvement in NDCG over the best-performing baseline GPT-3.5t (He et al., 2025).
@@ 2.2. Proposed Framework
We propose a framework to improve CRS as illustrated in Figure 1. Unlike previous models focusing on conversational contexts (Li et al., 2018; Chen et al., 2019; Zhou et al., 2020; Wang et al., 2022; He et al., 2023; Deng et al., 2021), we introduce dataset links between CRS and CF modalities to integrate user interaction information from both textual content and item representations generated via implicit feedback. Additionally, we incorporate LLMs in a zero-shot setting (He et al., 2023; Friedman et al., 2023) leveraging item representations generated by a pretrained CF recommender as supplementary information.
@@ 3.1. Results and Analysis
We assess performance of our framework against eight baselines as shown in Table 2. Our method consistently outperforms all baselines, demonstrating the effectiveness of integrating CF signals into conversational recommendations. Specifically, our framework achieves a 12.32% improvement in H@5 over the state-of-the-art baseline (He et al., 2025), GPT-3.5-t. Similarly, N@5 improves by 9.91%, further confirming the ranking quality of our method.
CF-based models perform well in leveraging historical user-item interactions but fall short in conversational settings due to their inability to incorporate contextual information. In contrast, CRS models such as ReDIAL and UniCRS incorporate conversational data but are limited in their ability to capture behavioral patterns. LLM-based CRS models leverage conversational context to achieve strong performance but tend to fail to generalize to less frequent items due to the absence of item representations. Our framework addresses these limitations by using CF embeddings with LLM-generated recommendations, resulting in superior accuracy across all scenarios. Unlike SASRec, which relies on historical interactions, or GPT-3.5-t, which lacks item-level embeddings, our framework bridges the gap between CF and conversational understanding.
@@ 4. Related Work
Exploiting collaborative signals can be done using matrix factorization (Koren et al., 2009) or a Graph Convolutional Network (GCN) (He et al., 2020). SASRec (Kang and McAuley, 2018) uses a self-attention network to exploit sequential patterns within collaborative signals. Using conversational context commonly involves exploiting the semantics of conversational text and recommending items in order to maximize alignment between conversation and item attributes. ReDial (Li et al., 2018) predicts user sentiment in the conversational context and feeds it to an autoencoder-based recommendation. EAR (Lei et al., 2020) extracts item attributes and attribute-specific user preferences from the conversational context and feeds them to an attribute-aware Bayesian Personalized Ranking (BPR) model for recommendation. KBRD (Chen et al., 2019) extracts item and non-item entities from the conversation and feeds them through a Relational Graph Convolutional Network (RGCN) to generate a knowledge-enhanced user preference representation. KGSF (Zhou et al., 2020) exploits a word-oriented knowledge graph via a GCN in addition to item-oriented knowledge to recommend items based on keywords in the conversation. UniCRS (Wang et al., 2022) jointly optimizes entity embeddings from an RGCN and contextual word embeddings from pretrained language models. Recent works have explored the direct use of conversational context by prompting LLMs (He et al., 2023; Friedman et al., 2023). ILM (Yang et al., 2024) uses item-text contrastive learning for learning text-aligned item representations from collaborative signals. (Zhang et al., 2023) proposes tuning a low-rank adapter associated with LLM and a collaborative information encoding module so that collaborative information can be exploited by LLMs. TALLRec (Bao et al., 2023) proposes instruction tuning of LLMs to improve their ability to exploit collaborative signals.

### v2

@@ Abstract.
journalyear: 2025copyright: ccconference: Companion Proceedings of the ACM Web Conference 2025; April 28-May 2, 2025; Sydney, NSW, Australiabooktitle: Companion Proceedings of the ACM Web Conference 2025 (WWW Companion ’25), April 28-May 2, 2025, Sydney, NSW, Australiadoi: 10.1145/3701716.3715486isbn: 979-8-4007-1331-6/25/04ccs: Information systems Recommender systemsccs: Computing methodologies Natural language generationccs: Computing methodologies Learning latent representations
@@ 1. Introduction
Conversational recommender systems (CRS) aim to elicit user preferences and generate recommendations through interactive dialogues (He et al., 2023; Li et al., 2018; Sun and Zhang, 2018; Wang et al., 2023). Unlike traditional recommenders, CRS leverages conversational context to adapt to user preferences. However, existing CRS approaches face two main challenges: (1) They do not leverage collaborative filtering (CF) signals with conversational context, leading to suboptimal recommendations (Zhou et al., 2020; Friedman et al., 2023), and (2) CRS datasets are typically sparse, making it difficult to develop robust item representations for accurate recommendations. To this end, a typical CRS consists of two components: (1) a generator to produce natural language responses and (2) a recommender to rank items aligned with user preferences (He et al., 2023; Li et al., 2018; Chen et al., 2019).
From a CRS model perspective, CF-based approaches excel at modeling user-item relationships but fail to leverage conversational context (Salem et al., 2014; Kang and McAuley, 2018). Similarly, knowledge-driven methods improve semantic understanding using external knowledge graphs but lack the collaborative signals needed to model item relationships (Zhou et al., 2020). Large language models (LLMs) show promise in CRS due to their natural language understanding capabilities (Brown et al., 2020; Chowdhery et al., 2023; Ouyang et al., 2022), but they rely at most on textual metadata, which constrains their ability to capture item relationships (He et al., 2025; Yang et al., 2024). Recent methods integrate CF embeddings with LLMs to improve recommendation tasks. For instance, A-LLMRec incorporates CF embeddings into prompts, while CoLLM maps them onto token embeddings (Kim et al., 2024; Zhang et al., 2023a). ILM aligns CF embeddings with textual embeddings but lacks conversational context (Yang et al., 2024). While these methods incorporate CF with LLMs to improve recommendations, they are not designed to utilize conversational context. Our framework bridges this gap by integrating CF signals and conversational data into a unified CRS approach.
@@
In this work, we make two key contributions: First, we construct a new Reddit-ML32M dataset by linking the Reddit-Movie dataset (He et al., 2023) with the MovieLens 32M dataset (Harper and Konstan, 2015). This dataset enriches the interaction space, increasing the total number of interactions from 51,148 to over 30 million and improving dataset density from 0.013% to 0.68%, as shown in Table 1. The dataset facilitates the development of richer item representations for CRS tasks by exploiting both conversational data with CF information at the item level. Second, we propose a framework that integrates conversational context from LLMs with CF-based item representations as depicted in Figure 1. GPT-3.5-turbo (OpenAI, 2024) [^1] We refer to this model as GPT-3.5t hereafter is used in a zero-shot setting as a backbone to generate initial recommendations from conversational context, which are then refined by ranking them against item embeddings generated by a pre-trained SASRec model (Kang and McAuley, 2018). This approach mitigates biases in LLM outputs, such as over-reliance on popular items or training data (He et al., 2023; Yang et al., 2024), while incorporating collaborative signals. Empirical evaluation across three paradigms: CF-based recommenders, CRS models, and LLM-based CRS without CF signals shows the effectiveness of our framework. Our approach achieves a 12.32% improvement in Hit Rate and a 9.9% improvement in NDCG over the best-performing baseline GPT-3.5t (He et al., 2025).
@@ 2.1. Problem Formulation
2.2. Dataset
We introduce Reddit-ML32M, a novel dataset that bridges Reddit Movie dataset for CRS tasks (He et al., 2023) with MovieLens 32M, a collaborative filtering dataset (Harper and Konstan, 2015). The key challenge is aligning items across these datasets. To address this, we use a two-phase linking method. First, IMDb IDs serve as unique identifiers, as they exist in both datasets. However, direct matching may fail when IMDb IDs are modified, reassigned, or deprecated over time, leading to inconsistencies between datasets. To resolve these cases, we use IMDbPY  [^3] https://github.com/MaximShidlovski23/imdbpy to retrieve historical IMDb IDs using exact title and release date matching, ensuring correct alignment even when an item’s IMDb ID has changed over time or been deprecated. This linking process enriches the interaction space, increasing total interactions from 51,148 to over 30 million, with a corresponding density improvement from 0.013% to 0.68% as shown in Table 1. We partition the dataset into 80% training, 10% validation, and 10% test sets.
@@ 2.3. Proposed Framework
We propose BridgeCRS, a framework designed to improve CRS, as illustrated in Figure 1 and outlined in Algorithm 1. Unlike previous models focusing on conversational contexts (Li et al., 2018; Chen et al., 2019; Zhou et al., 2020; Wang et al., 2022; He et al., 2023; Deng et al., 2021), we introduce dataset links between CRS and CF modalities to integrate user interaction information from both textual content and item representations generated via implicit feedback. Additionally, we incorporate LLMs in a zero-shot setting (He et al., 2023; Friedman et al., 2023) leveraging item representations generated by a pretrained CF recommender as supplementary information.
@@
Algorithm 1 BridgeCRS Algorithm
1:Conversation history $C$; pre-trained LLM $\mathcal{M}$; CF embedding matrix $E$; Candidate item set $\mathit{V}$
2:Ranked recommendation list $\hat{\mathit{I}}^{k}$
3:LLM-Based Recommendation
4:$\texttt{response}\leftarrow\mathcal{M}(\mathit{F}(\texttt{[TASK]},\texttt{[FORMAT]},C))$
5:Extract $\mathit{I}_{LLM}$ from response
6:Post-LLM Item Retrieval
7:$\mathit{I}_{LLM}\leftarrow\mathit{I}_{LLM}\cap\mathit{V}$triangleright Keep only items in candidate set
8:Embedding-Based Item Ranking
9:Retrieve embeddings $E_{LLM}\in\mathbb{R}^{m\times d}$ for $\mathit{I}_{LLM}$
10:Compute $s(j)=\max_{i\in\mathit{I}_{LLM}}\frac{\mathbf{e}_{i}\cdot\mathbf{e}_{j}}{\|\mathbf{e}_{i}\|\|\mathbf{e}_{j}\|},\quad\forall j\in\mathit{V}$
11:return $\hat{\mathit{I}}^{k}=\text{argsort}_{j\in\mathit{V}}(-s(j))$
@@ 3.1. Results and Analysis
We assess performance of our framework against eight baselines as shown in Table 2. Our method consistently outperforms all baselines, demonstrating the effectiveness of integrating CF signals into conversational recommendations. Specifically, our framework achieves a 12.32% improvement in H@5 over the state-of-the-art baseline (He et al., 2025), GPT-3.5-t. Similarly, N@5 improves by 9.91%, further confirming the ranking quality of our method. Note that H@1 and N@1 are identical because NDCG@1 reduces to Hit@1 due to the absence of rank discounting.
CF-based models perform well in leveraging historical user-item interactions but fall short in conversational settings due to their inability to incorporate contextual information. In contrast, CRS models such as ReDIAL and UniCRS incorporate conversational data but are limited in their ability to capture behavioral patterns. LLM-based CRS models leverage conversational context effectively but struggle with less frequent items due to the absence of item representations. Our framework addresses these limitations by using CF embeddings with LLM-generated recommendations, resulting in superior accuracy across all scenarios. Unlike SASRec, which relies on historical interactions, or GPT-3.5-t, which lacks item-level embeddings, our framework bridges the gap between CF and conversational understanding.
Generalization to Other Domains
The proposed dataset linking method is broadly applicable across domains where structured identifiers facilitate dataset alignment, such as ASINs in e-commerce and ISBNs in books (Ni et al., 2019; Wan and McAuley, 2018). This approach extends to industry settings where proprietary user interactions contain unique product identifiers, allowing collaborative item representations to be generated from implicit feedback sources such as purchase history and browsing patterns (Yoon et al., 2024). By leveraging structured identifiers and user interaction data, this framework can integrate collaborative signals into CRS tasks, which bridging structured and conversational modalities across diverse domains.
@@ 4. Related Work
Exploiting collaborative signals can be achieved using matrix factorization (Koren et al., 2009) or a Graph Convolutional Network (GCN) (He et al., 2020). SASRec (Kang and McAuley, 2018) uses a self-attention network to exploit sequential patterns within collaborative signals. Using conversational context typically involves leveraging the semantics of conversational text and recommending items to maximize alignment between conversation and item attributes. ReDial (Li et al., 2018) predicts user sentiment in the conversational context and feeds it to an autoencoder-based recommendation. The autoencoder-based recommender is similar to AutoRec (Sedhain et al., 2015) which is trained to produce unobserved user ratings from the observed user ratings. EAR (Lei et al., 2020) extracts item attributes and attribute-specific user preferences from the dialogue context and feeds them to an attribute-aware Bayesian Personalized Ranking (BPR) model for recommendation. KBRD (Chen et al., 2019) extracts item and non-item entities from the conversation and feeds them through a Relational Graph Convolutional Network (RGCN) (Schlichtkrull et al., 2018) to generate a knowledge-enhanced user preference representation. The RGCN is trained with an item-oriented knowledge graph from DBpedia (Lehmann et al., 2015). KGSF (Zhou et al., 2020) exploits a word-oriented knowledge graph via a GCN in addition to item-oriented knowledge to recommend items based on keywords in the conversation. The GCN associated with the word-oriented knowledge graph is trained on ConceptNet (Speer et al., 2017). UniCRS (Wang et al., 2022) jointly optimizes entity embeddings from an RGCN and contextual word embeddings from pretrained language models.
Recent works have explored the direct use of conversational context by prompting LLMs (Friedman et al., 2023; Hou et al., 2024). He et al. (He et al., 2023) feed dialogue context directly to LLMs for generating recommendations and do not use user-item interaction data. Liu et al. (Liu et al., 2023) demonstrate the use of LLMs for several recommendation tasks such as direct recommendation, rating prediction, sequential recommendation, explanation generation, and review summarization by constructing task-specific prompts. However, this approach performs poorly on direct and sequential recommendation (Zhang et al., 2023b). Gao et al. (Gao et al., 2023) propose adding user-item interaction history to the LLM prompts in addition to the dialogue context. However, LLMs cannot exploit complex user-item interactions from their textual representations. ILM (Yang et al., 2024) uses item-text contrastive learning for learning text-aligned item representations from collaborative signals. They train an item encoder to produce text-aligned item representations so that a frozen LLM may exploit item representations with existing parametric knowledge. Zhang et al. (Zhang et al., 2023a) propose encoding user-item interaction using traditional collaborative models and tuning a low-rank adapter associated with LLM so that collaborative information can be exploited by LLMs. TALLRec (Bao et al., 2023) proposes instruction tuning of LLMs to improve their ability to exploit collaborative signals.